Operations Report – February 3rd

Sol 6

Prepared by: Zac Trolley

Non-Nominal Systems: Greenhab temperature sensor, Suit 2, electrical systems.

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.7h
Solar: SOC 84%
Diesel: 76%
Propane: 61 psi
Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons
Water (static): 200 Gallons
Trailer to Static Pump used: No
Water (loft) – Static to Loft Pump used: Yes
Meter 130708.6

Toilet tank emptied: Yes

ATVs Used: None

Oil Added: None

ATV Fuel Used: None
# Hours the ATVs were Used today: 0
Notes on ATVs: none

Deimos rover used: Yes
Hours: Hours: 108.1
Beginning charge: 100%
Ending charge: 98%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 18.0h
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

Opportunity rover used: no
Hours: 13.5
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Curiosity rover used: Yes
Hours: 11.3h
Beginning charge: 100%
Ending charge: 92%
Currently charging: Yes

HabCar: Not used

General notes and comments: We are rotating the Rovers with the one plugin we have, and that should be sufficient if we are taking 2 rovers at a time per EVA.

It would seem that the electrical generation isn’t being regulated that well. The diesel and solar power systems will vary between 120V and 140V. No known reason why this is happening.

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crew’s ability to monitor the CapCom emails

Summary of suits and radios: There is a bent clasp on suit #1 on the helmet collar. It was probably bent during the EVA, the system is still functional.

Suit 2’s air supply is no longer functioning. The diagnostics we performed suggest that the switch isn’t operating properly. We are shelving the EVA suit.

Summary of Hab operations: Nominal

Summary of EVA operations: We are still brainstorming ways to combat the helmet fog. Without proper ventilation under the helmet ring, the facemasks rapidly fog up. It is becoming a safety issue on EVA. The other suit configuration seems to have better airflow.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: Nominal

Copyright © The Mars Society. All rights reserved. | Main Site